OpenStudy (calstu):

Can you solve this without L'Hopital's Rule: sqrt(x^2 + 8x) -x

satellite73 (satellite73):

you can't solve it with lhopital, it is not a fraction

jhonyy9 (jhonyy9):

@CalStu this is very easy to solve it simple just make equal zero ,add to both sides x ,square both sides ,and will get x^2 +8x = x^2 subtract from both sides x^2 and will get 8x = 0 divide both sides by 8 and will get x = 0 hope helped

OpenStudy (holsteremission):

I assume the limit is at either \(\pm \infty\)... \[\begin{align*} \lim_{x\to \pm\infty}\left(\sqrt{x^2+8x}-x\right)&=\lim_{x\to \pm\infty}\left(\sqrt{x^2+8x}-x\right)\times\frac{\sqrt{x^2+8x}+x}{\sqrt{x^2+8x}+x}\\[1ex] &=\lim_{x\to\pm\infty}\frac{\left(\sqrt{x^2+8x}\right)^2-x^2}{\sqrt{x^2+8x}+x}\\[1ex] &=\lim_{x\to\pm\infty}\frac{x^2+8x-x^2}{\sqrt{x^2+8x}+x}\\[1ex] &=\lim_{x\to\pm\infty}\frac{8x}{\sqrt{x^2}\sqrt{1+\frac{8}{x}}+x}\\[1ex] &=\lim_{x\to\pm\infty}\frac{8x}{|x|\sqrt{1+\frac{8}{x}}+x}\\[1ex] &=\lim_{x\to\pm\infty}\frac{8}{\frac{|x|}{x}\sqrt{1+\frac{8}{x}}+1} \end{align*}\]The limit then depends on the direction in which \(x\) is heading. You can use the definition of absolute value to finish the problem: \[|x|=\begin{cases}x&\text{if }x\ge0\\-x&\text{if }x<0\end{cases}\]